Suppose u(x) = cot x is an Integrating factor of the differential equation (sin x)? y" = 2y, find the general solution.


1
Expert's answer
2021-03-31T13:24:34-0400

Ans:- According to the question

integrating factor is

"\Rightarrow" "I.F.=e^{\int{p}dx}=Cotx"

"\Rightarrow" "\int{pdx}=ln{Cotx}"

"\Rightarrow" "p=-2\times Sec2x"

So then we arrange our question with respect to Integrating factor which has been given

As we see the question ? mark is coming "(sin x)? y" = 2y" then instead of ? "2cosx" written and "y''" written as "y'" .

Thus the question is "2{sin x}\times{cosx} y'= 2y"

"\Rightarrow y'-2y\times{Sec2x}=0"

Integrating factor is "Cotx"

then "y\times{Cotx}=\int{Cotx}\times{0}+c"

"\Rightarrow y\times{Cotx}=c" where c is an arbitrary constant